OpenStudy (anonymous):

What is √(−5) in the form a+bi?

OpenStudy (anonymous):

if you factor and split up the square root, it should be apparent... \[\sqrt{-5}=\sqrt{-1\cdot 5}=\sqrt{-1}\cdot \sqrt{5}=0+\sqrt{-1}\cdot \sqrt{5}=?\]

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

thank you